OpenStudy (anonymous):

Please check my answer! Perform the indicated operation and simplify. Come to the question so it can be displayed as an equation.

OpenStudy (anonymous):

\[((x^2+12x+36)/(x^2-8x+12))\div(x^2-36)\]

OpenStudy (anonymous):

My answer is probably wrong: (x+6)/(x-6)^2(x-2)

OpenStudy (mertsj):

\[\frac{(x+6)(x+6)}{(x-6)(x-2)} \times\frac{1}{(x-6)(x+6)}\]

OpenStudy (dape):

That is correct.

OpenStudy (anonymous):

Thank you! (:

OpenStudy (mertsj):

yw

OpenStudy (dape):

You're welcome =)